OpenStudy (anonymous):

Approximate the real number solution(s) to the polynomial function f(x) = x^3 + 2x^2 – 5x – 6 ? Help Please? Tried ..didnt get right.

OpenStudy (anonymous):

did you try cheating?

OpenStudy (anonymous):

Put x = -1 Will you get 0??

OpenStudy (anonymous):

are there 3 real sols

OpenStudy (anonymous):

-1 is one of them..

OpenStudy (anonymous):

No? I tried plugging the solutions in the equation but I dont think thats how you do It so i wanted someone to explain to me. Dont have to be a jerk.

OpenStudy (anonymous):

see I tried -1 , I guess I did it wrong becasue it didnt come out to equaling 0.

OpenStudy (anonymous):

this one factors, but it may be hard to see that it was cooked up try 1 first, if that doesn't work, try \(-1\) the only integer possibilities are \(\pm1,\pm2,\pm3,\pm6\)

OpenStudy (anonymous):

The three solutions will be: -1, -3 and 2..

OpenStudy (anonymous):

yes 3 solutions since the leading coefficient has a exponent of 3

OpenStudy (anonymous):

x^3 + 2x^2 – 5x – 6=0 x^2(x-2) +4x(x-2) +3=0 (x-2)(x^2 +4x+3)=0 (x-2)(x+1)(x+3)=0 x=-3,-1,2

OpenStudy (valpey):

I must have missed the part where we are asked to solve for f(x) = 0.

OpenStudy (anonymous):

Yes.. All the 3 solutions are real..